OpenStudy (anonymous):

In the reaction Mg (s) + 2HCl (aq) H2 (g) + MgCl2 (aq), how many moles of hydrogen gas will be produced from 75.0 milliliters of a 1.0 M HCl in an excess of Mg

OpenStudy (abmon98):

Number of moles=Concentration(mole/L)*Volume(L) There are 1,000 milliliters in a liter 1000 mL = 1 L 75.0 ml= x L Cross multiplication to find x L x=0.075 L Number of moles=0.075*1=0.075 moles 2 moles of HCl react with 1 mole of Mg to produce 1 mole of H2 and MgCl2 according to the balanced chemical equation. 1:2:1:1 Mg:H2 2:1 0.075: y moles of H2 0.075*1/2=0.0375